OpenStudy (anonymous):

Suppose a population of 50 crickets doubles in size every 4 months. How many crickets will there be after 5 years?

OpenStudy (anonymous):

125

OpenStudy (jim766):

y = 50(2)^15 =

OpenStudy (jim766):

50 is what you start with, 2 is how much the amount it multiplied 15 is the number of 4 month periods in 5 years. y = a(b)^x

OpenStudy (jim766):

much bigger than 125,,,

OpenStudy (anonymous):

oh i sorry i miss read years for months